If two planes leave the same airport at 1:00 PM, how many miles apart will

they be at 3:00 PM if one travels directly north at 150 mph and the other travels directly west at 200 mph?


Answer by checkley77(12569)

(2*150)^2+(2*200)^2=c^2
300^2+400^2=c^2
90,000+160,000=c^2
250,000=c^2
c=sqrt250,000
c=500 answer for the distance between the 2 planes 2 hours after they leave.